struct vehicle {

long carId;

short wheels:3;

short fuelTank : 6;

short weight;

}

what is the size of each one
Answer #1

long carld = 64 bits

short wheels : 3 = 16*3 = 48 bits

short fueltank : 6 = 16*6 = 96 bits

short weight = 16 bits